The vertices of a triangle are P(-6,1), Q(-2,-5) and R(8,1).
Find the equation of the perpendicular bisector of the side QR.

Find the slope of the median of the triangle that passes through point R.

Find the slope of the altitude of the triangle that passes through point Q.

Answer by mananth(16946)


The vertices of a triangle are P(-6,1), Q(-2,-5) and R(8,1).
Find the equation of the perpendicular bisector of the side QR.
slope formula
(y2-y1)/(x2-x1) Q(-2,-5) and R(8,1).
slope of QR = (1-(-5))/(8-(-2)) = 3/5
T a perpendicular line will have a slope -5/3 (negative reciprocal)
slope = -5/3 and passing through P (-6,-1)
Plug value of the slope and point ( -6 , -1 ) in
Y = m x + b
-1.00 = 10 + b
b= -1.00 - 10
b= -11
So the equation will be
Y = -5/3 x -11

Find the slope of the median of the triangle that passes through point R.
find the mid point of (PQ) (-4,-2)
and use slope form
Find the slope of the altitude of the triangle that passes through point Q.
This is done similarly